A solution is made by combining 10.0 mL of 18.0 M acetic acid with 5.58 g of sodium acetate and diluting to a total volume of 1.50 L.

Calculate the pH of the solution.

Express your answer using two decimal places.
Answer #1

first calculate the molarity of both solutions

M1V1 = M2V2

18.0 x 10 = M2 x 1500

M2 = 0.12 M

[CH3COOH] = 0.12 M

[CH3COONa] = (5.58 / 82.03)(1 / 1.50) = 0.045 M

now

pH = pKa + log [sodium acetate / acetic acid]

pKa of CH3COOH = 4.74 standard value

pH = 4.74 + log [0.045] / [0.12]

pH = 4.31